11. A neutral water molecule (H20) in its vapor state has an electric dipole moment of magnitude 6.2 x10 30 cm. How far apart are the molecules center of positive and negative charge?
Answer #1

electric dipole moment = charge * seperation

p = q*d

p = (10e)*d

6.2*10^-30 = 10*1.6*10^-19 * d

d = 3.875*10^-12 m

= 3.9 pm

so option C) 3.9 pm is the correct answer